OpenStudy (anonymous):

PLEASE HELP! A ramp is being used to unload the back of a truck. If the ramp is 5 feet long, and the back of the truck is 3 feet tall, how far will the end of the ramp be from the back of the truck? 2 ft 8 ft 5.8 ft 4 ft

OpenStudy (pfenn1):

Use the Pythagorean Theorem to solve.\[x^2 +y^2=z^2\]where x and y are sides of the triangle and z is the hypotenuse.

OpenStudy (anonymous):

is it 5.8?

OpenStudy (pfenn1):

No. How did you get that?

OpenStudy (anonymous):

I have no cule lol !is it 4?

OpenStudy (pfenn1):

Yes. It is 4.

OpenStudy (anonymous):

My firend is trying to help me and that is what she got!lol the 4 is my answer!

OpenStudy (pfenn1):

The equation you would have used is \[x^2+3^2=5^2\]Rearrange and solve for x\[x^2=25-9\]\[x^2=16;x=4\]

OpenStudy (anonymous):

that makes more sence than what me firend was trying to explain to me!lol